OpenStudy (anonymous):

The sum of two numbers is 14 and their difference is 6, write system of linear equations and solve the equation

OpenStudy (anonymous):

let x and y be the numbers sought: The sum is 14------->x+y=14 The difference is 6---->x-y=6 Substract 2nd equation from the first and get: x-x+y-(-y)=14-6----> 2y=8--->y=4 then replace y in any of the two equations and get x=14-4=10 or x=6+4=10 Your numbers are x=10 and y=4

OpenStudy (anonymous):

how cani graph it

OpenStudy (anonymous):

First equation is a line that intercepts x-axis in x=14 and y-axis in y=14. Draw it Second equation is a line that intercepts x-axis in x=6 and y-axis in y=-6. Draw it The cross point of the two linesis the solution of the equation

OpenStudy (anonymous):

check it!

OpenStudy (anonymous):

@amydavis82 have you done it already?

OpenStudy (anonymous):

yes thank you